# Discontinuous Galerkin discretization of conservative dynamical low-rank approximation schemes for the Vlasov–Poisson equation

### v1

@@ Abstract
A numerical dynamical low-rank approximation (DLRA) scheme for the solution of the Vlasov–Poisson equation is presented. Based on the formulation of the DLRA equations as Friedrichs’ systems in a continuous setting, it combines recently proposed conservative DLRA methods with a discontinuous Galerkin discretization. The resulting scheme is shown to ensure mass and momentum conservation at the discrete level. In addition, a new formulation of the conservative integrator is proposed which facilitates a projector splitting integrator. Numerical experiments validate our approach in one- and two-dimensional simulations of Landau damping. As a demonstration of feasibility, it is also shown that the rank-adaptive unconventional integrator can be combined with mesh adaptivity.
@@ 1 Introduction
As for DLRA methods, a recent research focus has been on the development of conservative methods, which is somewhat non-trivial to achieve. The ideas developed in [7] allowed for significant progress in this direction. By combining them with the rank-adaptive unconventional integrator from [4], also called basis update and Galerkin (BUG) integrator [3], the recent work [10] presents a robust numerical time stepping scheme for the factors $X_{i}$, $S_{ij}$ and $V_{j}$ in the DLRA representation (3) which is mass and momentum conservative and allows the rank $r$ to be chosen adaptively.
@@
The goal of the present work is to combine the conservative DLRA integrator from [10] with a discontinuous Galerkin (DG) discretization. This is motivated by the fact that the governing equations for the DLRA factors in the continuous setting take the form of Friedrichs’ systems, as worked out in [21]. DG methods possess favorable numerical properties for solving such systems, and, in addition, offer the possibility for adaptive mesh refinement. We make two main contributions. First, for a fixed discretization, we obtain a rank-adaptive and conservative DG-based DLRA approximation scheme for the Vlasov–Poisson equation. Second, we demonstrate how our approach of first formulating the DLRA scheme in a continuous setting and then discretizing the subproblems allows us to exploit the capabilities of DG solvers regarding mesh adaptivity. Compared to, e.g., [15], where the DLRA approach is applied to a full DG discretization of a linear kinetic equation, this approach offers much more flexibility. While the mesh adaptive version of our solver in its current form does not yet yield a conservative scheme due to the coarsening steps, we believe this approach to be a promising route for further development.
@@
In more detail, the outline and contributions of the paper can be summarized as follows. In section 2 we derive the equations of motion for the DLRA approach. We review the standard DLRA integrators in section 2.1 and the required modifications for the conservative integrator as proposed in [7, 10] in section 2.2. We pay special attention to formulating the equations in a continuous setting as Friedrichs’ systems in weak form, which makes it easier to enable spatially adaptive DG discretizations later. Our derivation of the equations for the conservative integrator in section 2.3 is based on a slightly different tangent space decomposition compared to [7, 10]. In particular, we present a new version of the modified L-step using a suitable block QR decomposition of the DLRA representation, which avoids forming normal equations. This L-step is then included in the rank-adaptive unconventional integrator, which is introduced in section 2.4. As a secondary result, our special QR representation also suggests a projector-splitting integrator allowing for fixed velocity basis functions which has not been featured in the original works. However, as we show in the numerical experiments, the projector splitting version is not conservative, and hence the unconventional integrator must to be used.
@@
Section 3 presents the main theoretical contribution. We introduce the DG discretization for the DLRA equations in sections 3.1 and 3.2. Special attention is paid to the approximation of derivatives and numerical fluxes at the element interfaces. As explained in [7, 10], the conservation properties of time integrators can be ensured by enforcing discrete counterparts of certain continuity equations arising in the Vlasov–Poisson system. In section 3.3, we rigorously prove these discrete continuity equations for mass and momentum conservation under the assumption that central fluxes are used in the DG discretization and the explicit Euler method is applied for time-stepping. The proof holds for arbitrary spatial dimensions. In section 3.4, we discuss spatial adaptivity in the DG discretization, though currently in a simplified form.
@@ 2 Conservative DLRA integrator
In this section, we recall the conservative DLRA integrator as first presented in [7] and further developed in [10]. With the goal of an adaptive DG discretization in mind, we will, however, derive the governing equations from a weak formulation of the Vlasov–Poisson equation as given in [11, Sec. 76.3]. Denoting the $L_{2}$ inner product on $\Omega=\Omega^{(x)}\times\Omega^{(v)}$ with
@@
which are related through
@@
Here, the functions $X_{1}(t),\dots,X_{r}(t)\in\mathcal{W}^{(x)}$ and $V_{1}(t),\dots,V_{r}(t)\in\mathcal{W}^{(v)}$ are usually assumed to be orthonormal systems for all $t$ without loss of generality, whereas this is not the case for the combined functions $K_{1}(t),\dots,K_{r}(t)$ and $L_{1}(t),\dots,L_{r}(t)$, respectively. Note that $K_{j}(t)\in\operatorname{span}\{X_{1}(t),\dots,X_{r}(t)\}$ and $L_{i}(t)\in\operatorname{span}\{V_{1}(t),\dots,V_{r}(t)\}$ for every $i$ and $j$.
@@ 2.1 Standard DLRA integrators
We first review the basic ideas of DLRA integrators before presenting the modifications proposed in [7, 10] for obtaining conservative integrators in the following subsections. Using the representations (5a), (5b) or (5c) at an initial time $t_{0}$ one can derive equations of motion separately for the matrix $S$, the functions $K_{1},\dots,K_{r}\in\mathcal{W}^{(x)}$, and the functions $L_{1},\dots,L_{r}\in\mathcal{W}^{(v)}$, respectively.
For example, given an initial function $f_{r}(t_{0})$ of the considered form, equations for the time evolution of $K_{1}(t),\dots,K_{r}(t)$ in (5b) on a time interval $[t_{0},t_{1}]$ are obtained by fixing the current $V_{j}(t_{0})$ for $j=1,\dots,r$ and then restricting the weak formulation (4) to the linear subspace
@@
Equations for the velocity functions $L_{1}(t),\dots,L_{r}(t)$ in the representation (5c) on a time interval $[t_{0},t_{1}]$ are obtained in an analogous way by fixing $X_{1}(t_{0}),\dots,X_{r}(t_{0})$ instead, and restricting the weak formulation (4) to the linear space
@@
    Perform a K-step on $T_{\bm{V}(t_{0})}$ by solving the Friedrichs’ systems (8) on $[t_{0},t_{1}]$ for new functions $K_{1}(t_{1}),\dots,K_{r}(t_{1})\in\mathcal{W}^{(x)}$. Choose an orthonormal basis $X_{1}(t_{1}),\dots,X_{r}(t_{1})$ of the new subspace $\operatorname{span}\{K_{1}(t_{1}),\dots,K_{r}(t_{1})\}$ with coefficients $\tilde{S}_{ij}$ such that $K_{j}(t_{1})=\sum_{i}\tilde{S}_{ij}X_{i}(t_{1})$.
@@
    Take $S(t_{0})=\tilde{S}$ as starting the point for a “backward” S-step on $T_{\bm{X}(t_{1}),\bm{V}(t_{0})}$, that is, solve the ODE (11) with $\partial S(t)-\ldots{}$ instead of $\partial S(t)+\ldots{}$ and using the already updated $X_{1}(t_{1}),\dots,X_{r}(t_{1})$ instead of $X_{1}(t_{0}),\dots,X_{r}(t_{0})$. This gives a new matrix $\hat{S}$ at time $t_{1}$.
@@
    Perform an L-step on $T_{\bm{X}(t_{1})}$ by solving the corresponding Friedrichs’ system for new functions $L_{1}(t_{1}),\dots,L_{r}(t_{1})\in\mathcal{W}^{(v)}$ with recomputed initial conditions $L_{i}(t_{0})=\sum_{j}\hat{S}_{ij}V_{j}(t_{0})$. Choose an orthonormal basis $V_{1}(t_{1}),\dots,V_{r}(t_{1})$ of $\operatorname{span}\{L_{1}(t_{1}),\dots,L_{r}(t_{1})\}$, and obtain coefficients $S_{ij}(t_{1})$ such that $L_{i}(t_{1})=\sum_{j}S_{ij}(t_{1})V_{j}(t_{1})$.
@@ 2.2 Conservative DLRA integrator
The exact solution to the Vlasov-Poisson equation possesses several physically relevant conservation properties such as mass, momentum and energy conservation. Ensuring this in DLRA simulations requires nontrivial modifications and has been addressed only recently by Einkemmer and Joseph in [7]. Consider for example the mass
@@ 2.3 Tangent space decomposition and modified projector splitting
As an extension of the ideas in [7, 10], in order to simplify the decomposition of the tangent space (16), we propose to consider a particular low-rank representation of a given $f_{r}\in\mathcal{M}_{r,\omega}$. Namely, after applying a suitable orthogonal transformation to the functions $K_{1},\dots,K_{r}$ using a variant of (block) QR decomposition, we can assume that the matrix $S$ in (13) is block triangular, that is,
@@
We further introduce the notation
@@
Then $f_{r}$ can be written as
$f_{r}=\omega\,\sum_{ab}S_{ab}X_{a}\otimes U_{b}+\omega\sum_{pb}S_{pb}Z_{p}\otimes U_{b}+\omega\sum_{pq}S_{pq}Z_{p}\otimes W_{q}.$ (18)
@@
$T_{\bm{Z},\omega}=\Big{\{}\,\omega\sum_{p}Z_{p}\otimes\widehat{\psi}_{p}^{(v)}:\widehat{\psi}_{p}^{(v)}\in\mathcal{W}^{(v)}_{\omega},\,(U_{a},\widehat{\psi}_{p}^{(v)})_{v,\omega}=0\,\Big{\}},$
@@
Based on the tangent space decomposition (19)–(20), we design a projector splitting integrator as well as a variant of the unconventional integrator. For this, we next derive the corresponding equations of motion as Friedrichs’ systems on subspaces $T_{\bm{V},\omega}$ (modified K-step), $T_{\bm{Z},\omega}$ (modified L-step) and $T_{\bm{Z},\bm{V},\omega}$ (modified S-step). Note that these spaces are defined solely by the basis sets $X_{1},\dots,X_{r}$ and $W_{m+1},\dots,W_{r}$. Hence, the assumption that $S$ is invertible is not essential for deriving DLRA integrators.
@@
$\widehat{L}_{\omega,p}(t_{0})=\omega\sum_{q}S_{pq}W_{q}\in\mathcal{W}^{(v)}.$
Recall that $(U_{a},\widehat{L}_{\omega,p}(t_{0}))_{v}=0$ for all $a$ and $p$. The goal is the time evolution of $\widehat{L}_{\omega,p}(t)$ under this orthogonality constraint which will yield
@@
(i.e. with the same “fixed” first part). To obtain $\widehat{L}_{\omega,p}(t_{1})$, the weak formulation (15) is restricted to the space $T_{\bm{Z},\omega}$, that is, to test functions of the form
$w=\omega\sum_{p}Z_{p}\otimes\widehat{\psi}_{p}^{(v)}.$
@@
Hence, (24) is an inhomogeneous Friedrichs’ system for $r-m$ functions $\widehat{L}_{\omega,p}(t)$ in $\mathcal{W}^{(v)}_{\omega}$ to be solved over the time interval $[t_{0},t_{1}]$. By construction, these will satisfy $(\widehat{L}_{\omega,p}(t),U_{a})_{v}=0$ for all $t\in[t_{0},t_{1}]$. We highlight that (24) represents an orthogonal projection of the system onto the space $T_{\bm{Z},\omega}$ and therefore differs from the formulation of the modified L-step in [10, Eq. (3)] based on normal equations.
@@
As already mentioned, the tangent space decomposition (19)–(20) suggests a projector splitting integrator based on the modified K-, S- and L-steps, which we present here for completeness. Although numerical experiments show that this scheme technically works (see section 4.1) it by itself does not yet achieve the desired conservation properties that motivated the introduction of the fixed functions $U_{a}$. For this purpose, the rank-adaptive unconventional integrator presented further below is necessary.
@@
    (Modified K-step) Denoting $\bm{V}^{0}=[V_{j}^{0}]_{j=1,\dots,r}$ and $\bm{K}^{0}=[K_{j}(t_{0})]_{j=1,\dots,r}$, solve the system (15) restricted to the subspace $T_{\bm{V}^{0},\omega}$ on the time interval $[t_{0},t_{1}]$ by solving the Friedrichs’ system (23) with initial condition $\bm{K}^{0}$. At time $t_{1}$, one obtains
@@
    (Modified S-step) Solve the system (15) restricted to the subspace $T_{\bm{Z}^{1},\bm{V}^{0},\omega}$ “backward” on the time interval $[t_{0},t_{1}]$ with initial condition $\tilde{f}$. This is achieved by solving (28) with $\partial_{t}\widehat{S}-\ldots$ (instead of $\partial_{t}\widehat{S}+\ldots$) on $[t_{0},t_{1}]$ with initial value $\tilde{S}$ from step 1 and corresponding matrices $\widehat{\mathsf{A}}^{(x,s)}$, $\bar{\mathsf{A}}^{(x,s)}$, $\widehat{\mathsf{B}}^{(x,s)}$, $\bar{\mathsf{B}}^{(x,s)}$ constructed from $\bm{X}^{1}$ and $\bm{V}^{0}$. At time $t_{1}$ one obtains
@@
    (Modified L-step) Solve the system (15) restricted to the subspace $T_{\bm{Z}^{1},\omega}$ on the time interval $[t_{0},t_{1}]$ with initial condition $\hat{f}$ by solving the modified L-step (24) with the data computed from $\hat{f}$. At time $t_{1}$ one obtains
    $\bar{f}_{r}=\omega\,\sum_{i}X_{i}^{1}\otimes L_{i}^{1}+\omega\,\sum_{p}Z_{p}^{1}\otimes\widehat{L}_{p}^{1}$

### v2

@@ Abstract
A numerical dynamical low-rank approximation (DLRA) scheme for the solution of the Vlasov–Poisson equation is presented. Based on the formulation of the DLRA equations as Friedrichs’ systems in a continuous setting, it combines recently proposed conservative DLRA methods with a discontinuous Galerkin discretization. The resulting scheme is shown to ensure mass and momentum conservation at the discrete level. In addition, a new formulation of the conservative integrator is proposed based on its interpretation as a tangent space projector splitting scheme. Numerical experiments validate our approach in one- and two-dimensional simulations of Landau damping. As a demonstration of feasibility, it is also shown that the rank-adaptive unconventional integrator can be combined with mesh adaptivity.
@@ 1 Introduction
As for DLRA methods, a recent research focus has been on the development of conservative methods, which is somewhat non-trivial to achieve. The ideas developed in [8] allowed for significant progress in this direction. By combining them with the rank-adaptive unconventional integrator from [5], also called basis update and Galerkin (BUG) integrator [3], in [12] a robust numerical time stepping scheme for the factors $X_{i}$, $S_{ij}$ and $V_{j}$ in the DLRA representation (3) has been presented, which is mass and momentum conservative and allows the rank $r$ to be chosen adaptively. Further improvements have been achieved recently in [10], where it is shown that augmentation of the spatial basis with gradients is not required for ensuring local conservation properties, hence leading to a more efficient algorithm.
@@
The goal of the present work is to combine the conservative DLRA integrator from [10] with a discontinuous Galerkin (DG) discretization. This is motivated by the fact that the governing equations for the DLRA factors in the continuous setting take the form of Friedrichs’ systems, as worked out in [23]. DG methods possess favorable numerical properties for solving such systems, and, in addition, offer the possibility for adaptive mesh refinement. A (non-adaptive) DG based DLRA algorithm for one spatial dimension has been worked out in [4]. With this work we make two main contributions. First, for a fixed discretization, we obtain a rank-adaptive and conservative DG-based DLRA approximation scheme for the Vlasov–Poisson equation. Second, we demonstrate how our approach of first formulating the DLRA scheme in a continuous setting and then discretizing the subproblems allows us to exploit the capabilities of DG solvers regarding mesh adaptivity. Compared to, e.g., [17], where the DLRA approach is applied to a full DG discretization of a linear kinetic equation, this approach offers much more flexibility. While the mesh adaptive version of our solver in its current form does not yet yield a conservative scheme due to the coarsening steps, we believe this approach to be a promising route for further development.
@@
In more detail, the outline and contributions of the paper can be summarized as follows. In section 2 we derive the equations of motion for the DLRA approach. We review the standard DLRA integrators in section 2.1 and the required modifications for the conservative integrator as proposed in [8, 12] in section 2.2. We pay special attention to formulating the equations in a continuous setting as Friedrichs’ systems in weak form, which makes it easier to enable spatially adaptive DG discretizations later. Our derivation of the equations for the conservative integrator in section 2.3 is based on a slightly different tangent space decomposition compared to [8, 12, 10]. In particular, we present a new version of the modified L-step using a suitable block QR decomposition of the DLRA representation, which avoids forming normal equations. This L-step is then included in the rank-adaptive unconventional integrator, which is introduced in section 2.4.
As a side result, our special QR representation also suggests a robust projector-splitting integrator allowing for fixed velocity basis functions which has not been featured in the original works. We present this alternative integrator briefly at the end of section 2.3. While it is not clear that this scheme is conservative by itself, it performs well in experiments and could be of some conceptual interest for further developments.
@@
Section 3 presents the main theoretical contribution. We introduce the DG discretization for the DLRA equations in sections 3.1 and 3.2. Special attention is paid to the approximation of derivatives and numerical fluxes at the element interfaces. As explained in [8, 12, 10], the conservation properties of time integrators can be ensured by enforcing discrete counterparts of certain continuity equations arising in the Vlasov–Poisson system. In section 3.3, we adopt the continuous approach from [10] to rigorously prove these discrete continuity equations for mass and momentum conservation in our setting under the assumption that central fluxes are used in the DG discretization and the explicit Euler method is applied for time-stepping. The proof holds for arbitrary spatial dimensions. In section 3.4, we discuss spatial adaptivity in the DG discretization, though currently in a simplified form.
@@ 2 Conservative DLRA integrator
In this section, we recall the conservative DLRA integrator as first presented in [8] and further developed in [12]. With the goal of an adaptive DG discretization in mind, we will, however, derive the governing equations from a weak formulation of the Vlasov–Poisson equation as given in [13, Sec. 76.3]. Denoting the $L_{2}$ inner product on $\Omega=\Omega^{(x)}\times\Omega^{(v)}$ with
@@
where the functions $X_{1}(t),\dots,X_{r}(t),K_{1}(t),\dots,K_{r}(t)\in\mathcal{W}^{(x)}$ and $V_{1}(t),\dots,V_{r}(t),L_{1}(t),\dots,L_{r}(t)\in\mathcal{W}^{(v)}$ are related through
@@
Clearly, $K_{j}(t)\in\operatorname{span}\{X_{1}(t),\dots,X_{r}(t)\}$ and $L_{i}(t)\in\operatorname{span}\{V_{1}(t),\dots,V_{r}(t)\}$ for every $i$ and $j$. The $X_{1}(t),\dots,X_{r}(t)$ and $V_{1}(t),\dots,V_{r}(t)$ are usually assumed to be orthonormal systems in $\mathcal{W}^{(x)}$ and $\mathcal{W}^{(v)}$ for all $t$ without loss of generality, whereas this is not the case for the combined functions $K_{1}(t),\dots,K_{r}(t)$ and $L_{1}(t),\dots,L_{r}(t)$, respectively.
@@ 2.1 Standard DLRA integrators
We first review the basic ideas of DLRA integrators before presenting the modifications proposed in [8, 12, 10] for obtaining conservative integrators in the following subsections. Using the representations (5a), (5b) or (5c) at an initial time $t_{0}$ one can derive equations of motion separately for the matrix $S$, the functions $K_{1}(t),\dots,K_{r}(t)\in\mathcal{W}^{(x)}$, and the functions $L_{1}(t),\dots,L_{r}(t)\in\mathcal{W}^{(v)}$, respectively.
For example, given an initial function $f_{r}(t_{0})$ of the considered form, equations for the time evolution of $K_{1}(t),\dots,K_{r}(t)$ in (5b) on a time interval $t_{0},t_{1}]$ are obtained by fixing the current $V_{j}(t_{0})$ for $j=1,\dots,r$ and then restricting the weak formulation (4) to the linear subspace
@@
Equations for the velocity functions $L_{1}(t),\dots,L_{r}(t)$ in the representation (5c) on a time interval $t_{0},t_{1}]$ are obtained in an analogous way by fixing $X_{1}(t_{0}),\dots,X_{r}(t_{0})$ instead, and restricting the weak formulation (4) to the linear space
@@
    Perform a K-step on $T_{\bm{V}(t_{0})}$ by solving the Friedrichs’ systems (7) on $t_{0},t_{1}]$ for new functions $K_{1}(t_{1}),\dots,K_{r}(t_{1})\in\mathcal{W}^{(x)}$. Choose an orthonormal basis $X_{1}(t_{1}),\dots,X_{r}(t_{1})$ for a subspace containing $K_{1}(t_{1}),\dots,K_{r}(t_{1})$ together with coefficients $\tilde{S}_{ij}$ such that $K_{j}(t_{1})=\sum_{i}\tilde{S}_{ij}X_{i}(t_{1})$.
@@
    Take $S(t_{0})=\tilde{S}$ as start point for a “backward” S-step on $T_{\bm{X}(t_{1}),\bm{V}(t_{0})}$, that is, solve the ODE (10) with $\partial S(t)$-$\ldots{}$ instead of $\partial S(t)+\ldots{}$ and using the already updated $X_{1}(t_{1}),\dots,X_{r}(t_{1})$ instead of $X_{1}(t_{0}),\dots,X_{r}(t_{0})$. This gives a new matrix $\hat{S}$ at time $t_{1}$.
@@
    Perform an L-step on $T_{\bm{X}(t_{1})}$ by solving the corresponding Friedrichs’ system for new functions $L_{1}(t_{1}),\dots,L_{r}(t_{1})\in\mathcal{W}^{(v)}$ with recomputed initial conditions $L_{i}(t_{0})=\sum_{j}\hat{S}_{ij}V_{j}(t_{0})$. Choose an orthonormal basis $V_{1}(t_{1}),\dots,V_{r}(t_{1})$ for a subspace containing of $L_{1}(t_{1}),\dots,L_{r}(t_{1})$ together with, and obtain coefficients $S_{ij}(t_{1})$ such that $L_{i}(t_{1})=\sum_{j}S_{ij}(t_{1})V_{j}(t_{1})$.
@@ 2.2 Conservative DLRA integrator
The exact solution to the Vlasov-Poisson equation conserves several physically relevant quantities such as mass, momentum and energy. Ensuring this in DLRA simulations requires nontrivial modifications and has been addressed only recently by Einkemmer and Joseph in [8]. Consider for example the mass
@@ 2.3 Tangent space decomposition and modified projector splitting
In [8] the equations of motion for $X_{i}$, $S$ and $V_{j}$ factors have been derived, but an inversion of $S$ was required for the updates of $V_{j}$. As an extension of their ideas, we derive a decomposition of the tangent space $T_{f_{r}}\mathcal{M}_{r,\omega}$ into orthogonal subspaces which allows to avoid the inversion of $S$, in the same spirit as in the original projector splitting integrator [21]. For that, consider a particular low-rank representation of a given $f_{r}\in\mathcal{M}_{r,\omega}$ tailored to the fact that the first $m$ velocity basis functions are fixed. We can assume that $f_{r}$ is initially given in the form $f_{r}=\omega\sum_{j}K_{j}\otimes V_{j}$. We we can then apply a suitable orthogonal transformation to the functions $K_{1},\dots,K_{r}$ using a variant of (block) QR decomposition, to obtain a new set of pairwise orthonormal functions $X_{1},\dots,X_{r}\in\mathcal{W}^{(x)}$ such that $K_{j}=\sum_{i}X_{i}S_{ij}$ as usual, but now with a block triangular matrix $S$ of the form
@@
Hence, without loss of generality, we can write $f_{r}$ as
$f_{r}=\omega\,\sum_{ab}S_{ab}X_{a}\otimes U_{b}+\omega\sum_{pb}S_{pb}Z_{p}\otimes U_{b}+\omega\sum_{pq}S_{pq}Z_{p}\otimes W_{q}$ (17)
where we introduced the notation
@@
$T_{\bm{Z},\omega}=\Big{\{}\,\omega\sum_{p}Z_{p}\otimes\widehat{\psi}_{p}^{(v)}:\widehat{\psi}_{p}^{(v)}\in\mathcal{W}^{(v)}_{\omega},\,(U_{a},\widehat{\psi}_{p}^{(v)})_{v,\omega}=0\,\Big{\}},$ (18)
@@
Based on the tangent space decomposition (19)–(20), we design a projector splitting integrator as well as a variant of the unconventional integrator. For this, we next derive the corresponding equations of motion as Friedrichs’ systems on subspaces $T_{\bm{V},\omega}$ (modified K-step), $T_{\bm{Z},\omega}$ (modified L-step) and $T_{\bm{Z},\bm{V},\omega}$ (modified S-step). Note that these spaces are defined solely by the basis sets $X_{1},\dots,X_{r}$ and $W_{m+1},\dots,W_{r}$ and hence can be defined even if $S$ is not invertible.
@@
$\widehat{L}_{\omega,p}(t_{0})=\omega L_{p}(t_{0})=\omega\sum_{q}S_{pq}W_{q}\in\mathcal{W}^{(v)}.$
We remark that due to our particular representation (17), this formula slightly differs from the corresponding formula in [12, Eq. (6)], where products of $S$ and its transpose are required for forming the new $L_{p}$ factors. Recall that $(U_{a},\widehat{L}_{\omega,p}(t_{0}))_{v}=0$ for all $a$ and $p$. The goal is the time evolution of $\widehat{L}_{\omega,p}(t)$ under this orthogonality constraint which will yield
@@
i.e. with the same “fixed” first part. The second sum in the above expression is a function in the subspace $T_{\bm{Z},\omega}$ of $T_{f_{r}(t_{0})}\mathcal{M}_{r,\omega}$ as defined in (18). Hence, equations for the time evolution of $\widehat{L}_{\omega,p}(t)$ are obtained by requiring the weak formulation (14) only for test functions from $T_{\bm{Z},\omega}$, that is, test functions of the form
$w=\omega\sum_{p}Z_{p}\otimes\widehat{\psi}_{p}^{(v)}$
@@
Hence, (24) is an inhomogeneous Friedrichs’ system for rmr-m functions $\widehat{L}_{\omega,p}(t)$ in $\mathcal{W}^{(v)}_{\omega}$ to be solved over the time interval $t_{0},t_{1}]$. By construction, these will satisfy $(\widehat{L}_{\omega,p}(t),U_{a})_{v}=0$ for all $t\in[t_{0},t_{1}]$. Eventually, we obtain $L_{p}(t_{1})=\omega^{-1}\,\hat{L}_{p}(t_{1})$. We highlight that (24) represents an orthogonal projection of the system onto the space $T_{\bm{Z},\omega}$ and therefore differs from the formulation of the modified L-step in [8, Eq. (11)] which requires the inversion of the matrix $S$.
@@
As already mentioned, the tangent space decomposition (19)–(20) suggests a projector splitting integrator based on the modified K-, S- and L-steps, which we present here for completeness. Although numerical experiments show that this scheme technically works (see section 4.1), keeping the functions $U_{a}$ in the bases by itself might not yet ensure the desired conservation properties. To guarantee this, the rank-adaptive unconventional integrator presented further below is necessary.
@@
    (Modified K-step) Denoting $\bm{V}^{0}=[V_{j}^{0}]_{j=1,\dots,r}$ and $\bm{K}^{0}=[K_{j}(t_{0})]_{j=1,\dots,r}$, solve the system (14) restricted to the subspace $T_{\bm{V}^{0},\omega}$ on the time interval $t_{0},t_{1}]$ by solving the Friedrichs’ system (23) with initial condition $\bm{K}^{0}$. At time $t_{1}$, one obtains
@@
    (Modified S-step) Solve the system (14) restricted to the subspace $T_{\bm{Z}^{1},\bm{V}^{0},\omega}$ “backward” on the time interval $t_{0},t_{1}]$ with initial condition $\tilde{f}$. This is achieved by solving (28) with $\partial_{t}\widehat{S}-\ldots$ (instead of $\partial_{t}\widehat{S}$$\ldots$) on $t_{0},t_{1}]$ with initial value $\tilde{S}$ from step 1 and corresponding matrices $\widehat{\mathsf{A}}^{(x,s)}$, $\bar{\mathsf{A}}^{(x,s)}$, $\widehat{\mathsf{B}}^{(x,s)}$, $\bar{\mathsf{B}}^{(x,s)}$ constructed from $\bm{X}^{1}$ and $\bm{V}^{0}$. At time $t_{1}$ one obtains
@@
    (Modified L-step) Solve the system (14) restricted to the subspace $T_{\bm{Z}^{1},\omega}$ on the time interval $t_{0},t_{1}]$ with initial condition $\hat{f}$ by solving the modified L-step (24) with the data computed from $\hat{f}$. At time $t_{1}$ one obtains
    $\bar{f}_{r}=\omega\,\sum_{i}X_{i}^{1}\otimes L_{i}^{1}+\omega\,\sum_{p}Z_{p}^{1}\otimes L_{p}^{1}$